Group Work 4, Section 12.1 Several Ways to Compute Double Integrals Consider the double integral SS (2 - x - y) dA, where R=a Complete this problem using a) lower left endpoints, and b) midpoints.
Answer #1

(nieren in hepeal ts Sanda Ciwan nectangle RTOJAT Hou f(x) (2x - 3) (9) OD Intants direction number of subintorals is m=2 InFayu No. a) J-1 f(0,5) = 15. f(15) L Sco- x - y) d4 = { { f(x ) DA - At Efro + $(15,9) + f(015) ++ (155) = (125) [2 +5+ is ]